OpenStudy (anonymous):

Will give medal!! whats 2+2?

OpenStudy (anonymous):

window.

OpenStudy (anonymous):

4

OpenStudy (aurorab):

really...?

OpenStudy (anonymous):

i almost used smiley's but tony's got it

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

0100 in binary

OpenStudy (anonymous):

2+2=0 , then 2=-2 give me the medal now!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

omg thats very hard lol

OpenStudy (anonymous):

Right :)

OpenStudy (anonymous):

are you forealz 2+2=4 or 2+2=0 2=-2 which is opposite